A jar contains 5- gram bolts & 10 gram bolts. The contents of the jar weigh 3.8 kg. If there are 460 bolts, how many 10 gram bolts are in the jar?
----------------------------------------------------------------------------
four more than 5 times a number is 24. Find the number.
Thanxs

Answer by stanbon(75887)

A jar contains 5- gram bolts & 10 gram bolts. The contents of the jar weigh 3.8 kg. If there are 460 bolts, how many 10 gram bolts are in the jar?
Let number of 10 gram bolts be "x"
Then number of 5 gram bolts is "460-x"
Weight of the 10 gram bolts is 10x grams
Weight of the 5 gram bols is 5(460-x)=2300-5x grams.
EQUATION:
wt. + wt.= 3800 grams
10x+2300-5x=3800
5x=1500
x=300 (number of 10 gram bolts)
460-x=160 (number of 5 gram bolts)
----------------------------------------------------------------------------
four more than 5 times a number is 24. Find the number.
Let the number be "x".
EQUATION:
5x+4=24
5x=20
x=4 (the number is "4"
Cheers,
Stan H.